OpenStudy (anonymous):

Alonzo is saving money to buy a game. So far he has saved $22, which is one-half of the total cost of the game. How much does the game cost?

OpenStudy (anonymous):

5/6*30=5*5=25

OpenStudy (anonymous):

@keepsecure93

OpenStudy (anonymous):

welcome to OS family.....